### v1

# OptiMind: Teaching LLMs to Think Like Optimization Experts

Abstract
Mathematical programming – the task of expressing operations and decision-making problems in precise mathematical language – is fundamental across domains, yet remains a skill-intensive process requiring operations research expertise. Recent advances in large language models for complex reasoning have spurred interest in automating this task, translating natural language into executable optimization models. Current approaches, however, achieve limited accuracy, hindered by scarce and noisy training data without leveraging domain knowledge. In this work, we systematically integrate optimization expertise to improve formulation accuracy for mixed-integer linear programming, a key family of mathematical programs. Our approach first cleans training data through class-based error analysis to explicitly prevent common mistakes within each optimization class. We then develop multi-turn inference strategies that guide LLMs with class-specific error summaries and solver feedback, enabling iterative refinement. Experiments across multiple base LLMs demonstrate that combining cleaned data with domain-informed prompting and feedback improves formulation accuracy by 14 percentage points on average, enabling further progress toward robust LLM-assisted optimization formulation.
1 Introduction
[…]
Mathematical optimization plays a critical role across many business sectors, from supply-chain management to energy systems to logistics planning, where effective decision-making relies on solving highly complex optimization problems. While practitioners can usually describe these problems in natural language, translating them into precise mathematical formulations that optimization solvers can process remains a skill-intensive bottleneck. Crafting a correct formulation requires precise definition of decision variables, objectives, and constraints, a skill that typically takes years of specialized training in operations research to develop.
Researchers and practitioners have begun exploring whether large language models (LLMs) can automate the formulation task: translating natural language problem descriptions into executable optimization models (Ramamonjison et al., 2022; Tang et al., 2024; Yang et al., 2025). Success on this front would lower a major barrier to broader use of optimization, democratizing its benefits. Yet current systems remain far from this goal. Accuracy is limited by training data quality, including lack of diversity and frequent errors from synthetic data generation. On the inference side, most existing approaches rely on a simple prompt that includes the natural language problem description and the modeling goal (e.g., generate an integer programming formulation in Pyomo, OR-Tools, or GurobiPy). This leaves out more structured prompting strategies that may yield stronger outcomes. The problem is further compounded by noisy benchmarks, such as ambiguous questions, missing data, and incorrect ground-truth values, where insufficient manual cleaning produces high error rates that obscure true performance. Critically, existing approaches underutilize domain expertise in benchmark quality control, training, and inference.
In this work, we study how optimization expertise can be systematically integrated – both at training time and inference time – to improve formulation accuracy at scale. Our first step is to clean existing training data. Motivated by our observation that existing LLMs often repeat similar mistakes within each optimization problem class, we first map problems into a fixed set of canonical classes (e.g., set cover, flow shop scheduling), then manually analyze a small subset of representative examples from each class to extract common formulation errors. We then use this error analysis to clean the entire training dataset by explicitly prompting strong LLMs to follow class-specific hints and avoid common error patterns, further using self-consistency to improve solution generation quality.
Beyond training data, we posit that achieving high accuracy in optimization formulation requires advanced inference techniques. Our error analyses and formulation hints derived from the training data are effectively incorporated into prompts to improve generation quality. Based on our observation that LLMs often generate noisy outputs, producing infeasible models or code with execution errors, we extend the inference pipeline into a multi-step process. In this pipeline, solver feedback – such as infeasibility detection, execution errors, and the previous optimal solution – is incorporated in conjunction with the hints to iteratively refine outputs, with self-consistency applied at each step to further stabilize performance. We henceforth refer to our training and inference framework as OptiMind, as it leverages human reasoning to improve a critical optimization task.
We evaluate our approach on the recently released gpt-oss-20b, applying both training-data and inference enhancements. To ensure reliable evaluation, we carefully re-clean three of the most challenging public benchmarks. On average, OptiMind improves performance by 14 percentage points across the evaluated benchmarks compared to the base model. An ablation study confirms that both training-data and inference components contribute significantly to these improvements. We further show that incorporating hints improves accuracy across multiple models, pointing to the robustness of our approach. Our contributions are summarized as follows:
  • We introduce a domain-informed error analysis framework that semi-automatically cleans and improves training data for optimization formulation tasks.
  • We propose a multi-turn inference pipeline that integrates class-specific error summaries with self-consistency and solver feedback to iteratively refine LLM outputs.
  • We conduct extensive empirical evaluation on three manually cleaned benchmarks, showing that our OptiMind framework – combining enhanced training data, domain expertise, and feedback-guided inference – substantially improves formulation accuracy achieving state-of-the-art results.
Together, these results demonstrate the importance of domain knowledge in making LLMs more reliable for optimization, advancing the broader goal of intelligent automation in decision-making. We plan to open-source our framework, data, cleaned benchmarks, and error analysis methods to enable further progress in the community.
Figure 1: OptiMind high-level overview.
Figure 2: From problem description to solution.
2 Description of the Formulation Task
The formulation task, which is the focus of this paper, consists in translating natural language problem descriptions into executable optimization models. To make this precise, we first describe the class of optimization models considered.
Mixed-Integer Linear Programming (MILPs). MILP is the class of optimization problems where some decision variables are constrained to be integers, and relationships among variables are linear in the form of constraints and an objective function. A MILP problem is formulated as $\min\{c^{\top}x:Ax\leq b,x_{j}\in\mathbb{Z}~\forall j\in I\},$ where $x\in\mathbb{R}^{n}$ is the vector of decision variables, $c\in\mathbb{R}^{n}$ is the cost vector, $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{R}^{m}$ define the linear constraints, and $I\subseteq\{1,\dots,n\}$ indicates the variables required to be integer-valued. MILPs are very general and used extensively to model complex decision-making problems involving both discrete choices and continuous variables, capturing applications in supply-chain optimization, scheduling, network design, resource allocation, and much more (Fleuren et al., 2013; Kroon et al., 2009; Durán et al., 2007; Lee et al., 2009).
The Formulation Task. The input of the task is a complete description in natural language of the decision-making problem that one wishes to model and solve, including all the required data. For example, in the context of manufacturing planning, this data typically includes demand values for the products on different periods, machine capacities or other resource quantities, etc. The desired output of the task is an MILP formulation of the input problem. Concretely, the output format we consider is a Python code that specifies the decision variables, constraints, and objective function of the output MILP; this leverages the strong ability of current LLM models of producing Python code. In addition to defining the MILP formulation, the output code also has commands to execute a solver (e.g., Gurobi (2025)) and a short routine to print the optimal decisions. Running the output code thus produces the complete set of decisions ready to be employed by the user. Fig. 2 illustrates this process, and in Appendix A.1 we provide a complete input/output example.
3 Related Work
A growing body of work focuses on building benchmarks for natural language to optimization formulation translation. NL4LP (Ramamonjison et al., 2022) was one of the early benchmarks introduced in a NL4OPT competition, focusing on Linear Programming (LP) problems. Mamo (Huang et al., 2024) expands the scope to MILPs, with Mamo Easy and Mamo Complex reflecting different difficulty levels. NLP4LP (AhmadiTeshnizi et al., 2024) contains LP and ILP problems collected from textbooks and lecture notes. Xiao et al. (2024) released ComplexOR, a benchmark of OR problems collected from both industrial and academic scenarios. IndustryOR (Tang et al., 2024) provides 100 real-world OR problems across eight industries, while OptiBench (Yang et al., 2025) extends coverage to nonlinear and tabular problems. OptMATH (Lu et al., 2025) further introduces GPT-synthesized benchmark with longer natural language contexts and complex constraints. A major limitation in current benchmarks is their high error rates. Both our own experience and a recent survey (Xiao et al., 2025) reveal frequent issues, with up to 30–50% error rates, including missing data, ambiguous formulations, and incorrect ground-truth answers. We thus have dedicated significant efforts to cleaning some of the benchmarks.
Several prompting strategies have been developed to improve formulation accuracy. Chain-of-experts (Xiao et al., 2024) and OptiMUS (AhmadiTeshnizi et al., 2024) adopt agentic frameworks that split the task into specialized LLM agents for modeling, programming, and evaluation. Search-based methods such as AutoFormulation (Astorga et al., 2025) use Monte-Carlo Tree Search to separately construct variables, constraints, and objectives, while pruning equivalent formulations at the symbolic level. More recently, OptiTree (Liu et al., 2025) further decomposes complex optimization tasks into simpler subproblems to improve the overall solution. Multiple works have explored fine-tuning strategies for optimization datasets. ORLM (Tang et al., 2024), ReSocratic (Yang et al., 2025), and OptMATH (Lu et al., 2025) apply supervised fine-tuning (SFT) on LLM-synthesized datasets to improve performance on IndustryOR, OptiBench, and OptMATH, respectively, while LLMOpt (Jiang et al., 2025) uses Kahneman-Tversky Optimization (KTO) on synthetic data. SIRL (Chen et al., 2025) combines SFT with RL to further boost results. However, reproducibility of these works remains a challenge: some models are highly prompt-sensitive and complete training data and latest checkpoints are often unavailable.
4 Method
We now detail the complete pipeline of our framework; see Fig. 3 for an overview.
Figure 3: An overview of our training data cleaning and multi-turn inference pipeline.
4.1 Test Data Cleaning
We consider three of the most challenging optimization formulation benchmarks – IndustryOR, Mamo Complex, and OptMATH – where even the strongest models at the time only report accuracies up to 20%–50% (Jiang et al., 2025; Liu et al., 2025; Lu et al., 2025). By contrast, earlier LP-centric benchmarks (e.g., NL4LP (Ramamonjison et al., 2022), NLP4LP (AhmadiTeshnizi et al., 2024)) already see 90–95%+ accuracy (see e.g. (Liu et al., 2025; Lu et al., 2025)), highlighting substantial headroom on these harder tasks. This raises a natural question: where do existing models fall short, and how can we improve?
On closer inspection, however, we find a surprising fact: many errors stem not from model capability but from issues in the benchmarks themselves: missing or ambiguous problem data, incorrect reference answers, and rounding inconsistencies in the evaluation pipeline, resulting in 30% - 60% test instances being incorrect. Despite recent efforts to clean these benchmarks (Tang et al., 2024; Chen et al., 2025), such issues remain widespread.
To ensure the reliability of evaluation, we carefully re-cleaned the test data of these three benchmarks. This was a non-trivial process, requiring over a month of manual effort from a team of optimization experts (with experience level of Professors and PhD students in Operations Research). The errors include missing data, ambiguous problem descriptions, integral vs. fractional variables and more. See Appendix A.9 for details, including how we corrected the issues. After cleaning, we observe a remarkable gain in the accuracy across all three benchmarks. With the same gpt-oss-20b model and inference strategy, the average accuracy increases from 40%–60% on the original releases to 70%–90% on our corrected sets. We will release our fixes and annotations for each problem to make results comparable across papers and better reflect true model ability.
4.2 Problem-Class Specific Error Analysis
While highly capable LLMs perform strongly on the cleaned benchmarks, in this work we take the ambitious goal of further improving their performance in optimization formulation. Upon examining their outputs on a small training subset, we find that formulation mistakes still occur, and often similar types of mistakes recur within the same optimization category: for example, in the Traveling Salesman Problem (TSP), the LLMs often mess up the subtour elimination constraint by incorrectly applying it to the fixed starting node (Fig. 4, middle box). Similarly, we see frequent mistakes in the signs of flow-conservation constraints in network-flow and inventory management problems; for example, if $B[u]$ denotes the net supply of a node $u$ (i.e., positive if there is extra supply), the flow conservation equation should read [outflow from $u$] - [inflow into $u$] =B[u]=B[u], but we often found the outflow and inflow terms swapped. This suggests a simple yet intuitive intervention: summarize short, targeted hints that capture the most common failure modes in each optimization category and attach the appropriate hint when solving problems from that category.
To this end, we develop a class-based error analysis. We classify all problems in the open-source optimization training sets (OptMATH (Lu et al., 2025) and OR-Instruct (Tang et al., 2024), which is the training data corresponding to the IndustryOR benchmark) into the 50 seed classes defined in the OptMATH problem generators; these classes provide a wide coverage of typical MILP-type problems. For each class, we run gpt-oss-20b to produce answers and select 10–20 instances where its answer disagrees with the original ground-truth label. This label mismatch can be attributed either to model errors in the solution generation or quality issues with the questions themselves (details in §4.3). Our optimization experts manually review each item to identify the source of the disagreement; when the model is at fault, we write a short error summary and a concrete hint that would have prevented the mistake (see Fig. 4 (left) for an example). We then aggregate these annotations into a dictionary that maps each problem type into a list of (error summary, preventive hint) pairs.
Hint Example for TSP Error description: "When solving TSP, it is easy to make mistakes by not enforcing the MTZ subtour elimination constraints correctly." Hint: "When solving TSP using MTZ, fix one city’s position (e.g., u[0]=0) and apply u[i] - u[j] + n*x[i,j] <= n - 1 for the *remaining* cities i, j != 0 with i != j to ensure no subtours are created; do *not* apply this constraint u[i] - u[j] + n*x[i,j] <= n - 1 for the city whose position you fixed (e.g., city 0)"
Response without Hint Thought: "We need to solve TSP with 6 nodes, directed distances given (not symmetric). We need to formulate TSP as integer program: binary variables x[i,j] = 1 if travel from i to j. Constraints: each city has exactly one outgoing edge and one incoming edge. Subtour elimination: use Miller-Tucker-Zemlin (MTZ) constraints: u[i] variables continuous between 1 and n. For i != 0, u[i] >= 2, <= n. For all i!=j, u[i] - u[j] + n * x[i,j] <= n-1. Or standard: u[i] - u[j] + n*x[i,j] <= n-1. Also need to ensure start city 0? Usually we set u[0] = 0. Or we can set u[0] = 0 and enforce others between 1 and n-1. But with 6 nodes, n=6. We’$l$ implement accordingly."
Response with Hint Thought: "We need to solve TSP with 6 nodes. Provide model: binary variables x[i,j] for i != j, objective minimize sum cost[i][j]*x[i,j]. Constraints: each node outdegree 1, indegree 1. Use MTZ subtour elimination: choose node 0 as fixed with u[0]=0, u[i] integer 1..n-1 for i>0. For all i != j, i,j > 0? Actually apply for all i != j, i>0 and j>0? The hint says not to apply constraint for city whose position fixed. So for i != j, i>0, j>0, i != j: u[i] - u[j] + n*x[i,j] <= n-1."
Figure 4: Sample hint for a problem type (Traveling Salesman Problem, TSP) and comparison of the model’s response before and after applying hints.
This outcome of the error analysis becomes a core component of both our training-data cleaning and our inference pipeline. Fig. 4 (right) demonstrates the power of a single hint in preventing a frequent subtour-related error. Fig. 7 (left) demonstrates the generalization capability of our hints: even without closely matched training items, the same hint family reduces mistakes on the Mamo-Complex benchmark on the biggest problem classes, yielding a 16.6% gain in overall accuracy.
4.3 Training Data Cleaning
Building on the error analysis above, we now examine how to leverage training data to systematically improve the base model. A common approach to improve LLMs for optimization formulation is by supervised fine-tuning on the training sets attached to each benchmark. However, we find that the training data from the datasets we consider (OR-Instruct and OptMATH) exhibits quality issues that mirror, and sometimes exceed, those in the test sets. First, many training solutions (reasoning, code, and final answers) were synthesized by older LLMs (e.g., OR-Instruct from Tang et al. (2024) uses gpt-4 and OptMATH from Lu et al. (2025) uses DeepSeek-V3), and we often observe low-quality or internally inconsistent outputs that would propagate errors into SFT. Second, many training questions contain missing parameters or ambiguous phrasing, akin to the issues we documented in the benchmarks. We exemplify this issue in Appendix A.8.
Unlike test benchmarks, these training corpora are large, so exhaustive manual relabeling is impractical. We therefore design a semi-automated cleaning pipeline that combines targeted expert review with scalable LLM-assisted checks. We pursue two complementary directions in order to obtain a higher-quality training dataset suitable for learning: (i) improve solution quality and labels, and (ii) improve question quality and clarity.
Improve solution quality. To improve “ground-truth” quality and balance across the datasets, we:
  • Balance classes. From the bigger dataset, OptMATH, we sample 100 training instances per problem class when available; for classes with fewer than 100 instances, we take them all.
  • Solution regeneration using error analysis and majority voting. Here we employ a simplified version of our inference process described in more detail in §4.5. That is, we use gpt-oss-20b augmented with the class-specific error summaries and hints described in §4.2 to reduce recurrent modeling mistakes. We use majority vote with K=64K{=}64 samples, yielding higher-quality solutions.
  • Filter unresolvable items. Finally, we drop problems where neither the original code nor the regeneration process produce a valid result.
This process yields $2700$ cleaned items for OR-Instruct and $2600$ for OptMATH. Of these, $602/2700$ in OR-Instruct and $577/2600$ in OptMATH have answers that differ from the original labels.
Improve question quality. To address the issue of missing data and ambiguous description:
  • Detect and fill missing parameters. We automate detection and completion of missing fields using the OpenAI o4-mini model, flagging $180/2700$ items in IndustryOR and $500/2600$ in OptMATH as incomplete and filling them with validated values. We then manually checked a few samples by our optimization experts to verify correctness.
  • Resolve ambiguity with expert edits. We manually clean classes with systematic ambiguity (e.g., Job Shop); an example is in Appendix A.8. In general, such ambiguity cannot be resolved reliably in a fully automated way; broader coverage of additional types is left to future work.
4.4 Supervised Fine-Tuning
We use supervised fine-tuning (SFT) to strengthen the model’s formulation and coding ability. Concretely, we fully fine-tune gpt-oss-20b on our cleaned training dataset. Let the SFT dataset be $\mathcal{D}_{\mathrm{SFT}}={(x_{i},y_{i})}_{i=1}^{N}$, where $x_{i}$ is the problem description and $y_{i}$ is the completion sequence formed by concatenating the model’s thinking tokens, the mathematical formulation, and the solver code. We train with standard sequence-to-sequence loss: $\mathcal{L}_{\mathrm{SFT}}(\theta)=-\mathbb{E}_{(x,y)\sim\mathcal{D}_{\mathrm{SFT}}}\sum_{t=1}^{|y|}\log p_{\theta}\bigl(y_{t}\mid y_{<t},x\bigr)$, where $\theta$ denotes all trainable parameters of the model; $p_{\theta}(\cdot)$ is the model’s output distribution; $y_{t}$ is the target token at position $t$; and $y_{<t}:=(y_{1},\ldots,y_{t-1})$ are the prefix of previously generated target tokens. Full training details are provided in the Appendix A.6.1.
4.5 Inference
We adopt a three-stage inference pipeline with (i) error-aware prompting, (ii) self-consistency via majority voting, and (iii) multi-turn correction with tool feedback. As shown in §5, these components reinforce each other to form a robust pipeline for solving optimization problems.
Improved prompts with error-analysis-based hints. We incorporate the class-based error analysis from §4.2 into the inference process. The model first classifies each test instance into one of the 50 classes that were defined in the training data error analysis, then augments its prompt with all error–hint pairs from that class to guide solution generation. Since the training instances from the same class can differ in their characteristics (e.g., some multi-period inventory problems allow backorders while others do not), we clarify in the prompt that hints should be applied only when relevant, allowing the model to ignore those hints that do not fit the problem description. This category-based targeted application helps the model avoid recurrent pitfalls and reduces common mistakes. On top of the hints we obtained from our analysis, we add general guidelines for correct formulation, which are derived from our optimization experts’ experience.
Self-consistency with majority voting. We sample multiple reasoning traces and take the self-consistent answer that appears most frequently. Taking the majority voting solution dampens stochastic variance from sampling, stabilizes final predictions, and yields strong gains even under otherwise simple prompts.
Multi-turn correction. We then run a multi-turn loop to further reduce errors and leverage the model’s self-improvement. After each round, we extract and execute the generated Python code, and collect system feedback containing Gurobi logs (e.g., the optimal solution found or whether infeasibility was detected) or execution errors (when code fails to run). We feed this feedback to the model, ask it to assess correctness, and, if needed, revise the code. With strong reasoning models such as gpt-oss, most low-level coding errors are fixed within a few turns, and the model can also self-correct modeling issues. Appendix A.5 provides an example of nontrivial self-corrections.
Putting it together, the full inference pipeline proceeds as follows: the initial turn classifies the problem; the next turn solves it using the returned hints for the predicted class(es). We generate $K$ solutions and select the majority-vote answer. We then return the solutions and system feedback to the model for self-correction, again aggregating by majority across multiple samples per turn. This correction loop is repeated for $M$ rounds. The pseudocode is provided in Alg. 1 in Appendix A.4.
5 Experiments
5.1 Experimental Setup
Datasets. Our evaluation is performed on our rigorously cleaned versions of three challenging and widely-used benchmarks: IndustryOR (Tang et al., 2024), Mamo-Complex (Huang et al., 2024), and OptMATH (Lu et al., 2025), each with around 100-160 problems (see Appendix A.9 for details). We selected these datasets as they are commonly used in previous works and represent some of the most complex formulation tasks in the literature, providing a strong signal for model capabilities. While we also considered other benchmarks such as OptiBench (Yang et al., 2025) (605 problems) and ComplexOR (Xiao et al., 2024) (18 problems), our initial experiments revealed that performance on these datasets has either saturated or does not effectively differentiate between models of varying scales; moreover, their sizes are either too small (yielding a noisy evaluation signal) or too large to feasibly clean with the same rigor. A more detailed discussion of our dataset selection and the comprehensive cleaning process is provided in Appendix A.8.
Metrics. Our primary metric is solution accuracy, which we report at the first generation turn (Turn 1) and after five turns of iterative self-correction (Turn 5). Our prompts require an executable Python script using GurobiPy to formulate and solve the problem. Following SIRL (Chen et al., 2025), a solution is correct if its objective matches the ground truth within both relative and absolute error of $10^{-6}$. To extract this value, we insert a print statement emitting the optimal objective with a unique tag and parse it from the execution log, as in SIRL. We also assess the effect of self-consistency by comparing results without majority voting ($K=1$) against results with majority voting over $K=8$ samples, grouping answers within a relative and absolute tolerance of $10^{-6}$ to account floating-point variations. To ensure statistical robustness, all reported results are averaged across 10 independent experiments using different random seeds.
Baselines. To comprehensively assess our contributions, we compare OptiMind against several baselines: (1) we compare with the gpt-oss-20b base model without any fine-tuning to quantify the gains from our training methodology; (2) at inference, we test a variant using only basic prompts without our class-specific error hints to verify the effectiveness of our domain-informed guidance; (3) we also benchmark our model against other frontier models, including GPT-o4-mini and GPT-5 and (4) an open-source models of similar size, Qwen3-32B; (5) furthermore, we re-evaluated SIRL-Gurobi-32B (Chen et al., 2025) for direct comparison with the current state-of-the-art models in the field of optimization. Note that we were unable to replicate the reported performance of other open-source models like OptMATH and LLMOpt.
5.2 Main Results
Figure 5: Average accuracy for different models (without majority voting). Our model outperforms the open-source models, and performs comparably to the larger GPT-5 model.
Table 1: Accuracies under different base models and hint settings on expert-cleaned benchmarks.
IndustryOR Mamo-Complex OptMATH
Model K Hints 1 turn 5 turns 1 turn 5 turns 1 turn 5 turns
gpt-5 1 no $78.57\pm 3.09$ $81.12\pm 1.79$ $87.86\pm 1.55$ $89.36\pm 1.92$ $90.15\pm 0.69$ $92.18\pm 0.78$
yes $83.50\pm 2.84$ $85.20\pm 1.25$ $92.57\pm 1.19$ $93.42\pm 0.62$ $89.21\pm 1.78$ $90.31\pm 1.04$
o4-mini 1 no $76.19\pm 1.05$ $82.14\pm 2.11$ $75.95\pm 2.38$ $87.78\pm 1.15$ $80.31\pm 2.02$ $87.18\pm 0.89$
yes $82.14\pm 2.47$ $5.88\pm 1.19 88$$8.09\pm 2.02$ $91.67\pm 1.52$ $84.21\pm 2.17$ $90.46\pm 1.16$
Qwen3-32B 1 no $64.28\pm 3.00$ $9.39\pm 3.00 66$$6.47\pm 1.64$ $77.14\pm 3.18$ $79.01\pm 2.41$ $79.49\pm 1.84$
yes $69.13\pm 2.00$ $74.23\pm 3.85$ $78.00\pm 1.73$ $84.76\pm 2.12$ $76.95\pm 1.35$ $80.85\pm 1.35$
SIRL-Gurobi32B 1 no $1.36\pm 1.39 55$$5.27\pm 1.63$ $64.00\pm 1.41$ $72.67\pm 1.37$ $70.93\pm 1.94$ $75.46\pm 1.62$
yes $49.38\pm 3.02$ $53.06\pm 3.61$ $75.61\pm 2.05$ $78.57\pm 0.95$ $74.06\pm 1.39$ $76.71\pm 2.30$
gpt-oss-20b 1 no $67.14\pm 3.46$ $73.67\pm 2.88$ $70.00\pm 1.93$ $86.57\pm 0.91$ $76.32\pm 2.63$ $86.32\pm 2.12$
yes $\bf{72.45\pm 1.66}$ $\bf{76.83\pm 1.52}$ $83.81\pm 3.05$ $89.90\pm 1.73$ $79.84\pm 2.04$ $86.56\pm 1.75$
Our SFT 1 no $68.36\pm 2.92$ $74.92\pm 2.42$ $74.42\pm 1.15$ $86.05\pm 1.55$ $80.35\pm 3.10$ $\bf{87.94\pm 1.41}$
yes $70.26\pm 1.90$ $74.20\pm 2.61$ $\bf{85.85\pm 0.76}$ $\bf{90.41\pm 0.58}$ $\bf{81.13\pm 1.59}$ $87.39\pm 1.05$
gpt-oss-20b 8 no $76.22\pm 3.15$ $78.37\pm 2.49$ $6.09\pm 3.55 88$$8.76\pm 1.72$ $86.95\pm 3.08$ $89.53\pm 2.27$
yes $\bf{80.10\pm 2.69}$ $79.69\pm 2.98$ $90.04\pm 2.76$ $90.72\pm 1.08$ $\bf{88.36\pm 2.06}$ $\bf{89.14\pm 1.40}$
Our SFT 8 no $75.85\pm 1.56$ $79.25\pm 0.58$ $7.61\pm 0.89 88$$8.19\pm 0.78$ $7.76\pm 5.31 88$$8.28\pm 4.00$
yes $78.91\pm 2.57$ $\bf{80.91\pm 1.77}$ $\bf{92.14\pm 0.61}$ $\bf{92.02\pm 0.71}$ $7.76\pm 1.19 88$$8.28\pm 2.06$
We now assess the impact of the components of OptiMind’s pipeline on solution accuracy. Accuracy improvements are reported as absolute percentage points.
Results with OptiMind’s training data cleaning. First, SFT training fed with the cleaned training data processed through OptiMind has demonstrable gains. Focusing on the gpt-oss-20b model and isolating training effects under the single-turn without hints, our SFT model outperforms the base model by $1.2\%$ on IndustryOR, $4.4\%$ on Mamo-Complex, and $4.0\%$ on OptMATH. Improvement gains remain under 5-turn self-correction setting (see Table 2 for details). Additional ablation study can be found in §5.3.
Table 2: GPT-OSS-20B vs. our SFT model under no-hint prompting, comparing single-turn ($K=1$) and 5-turn self-correction ($K=1$). We report SFT improvements over the base in percentage points.
Benchmark Base (1 turn) SFT (1 turn) Gain (%) Base (5 turns) SFT (5 turns) Gain
IndustryOR 67.14 68.36 +1.22 73.67 74.92 +1.25
Mamo-Complex 70.00 74.42 +4.42 86.57 86.05 -0.52
OptMATH 76.32 80.35 +4.03 86.32 87.94 +1.62
Results with OptiMind’s error analysis. Additionally, using class-specific error analysis and the associated hints at inference consistently lift single-turn accuracy across models and datasets. Typical gains from “no hint” to “with hints” across the baselines and our SFT model (single-turn, no majority voting) range from are $3\%$ to $6\%$ on IndustryOR, $11\%$ to $14\%$ on Mamo-Complex, and $1\%$ to $4\%$ on OptMATH, with rare small dips around $1\%$ to $2\%$ on OptMATH for a few models. Interestingly, even GPT-5 sees a significant improvement in accuracy when using the hints (up to $4.93\%$ on IndustryOR), suggesting that hints can enhance the performance of even very strong models, encoding domain-specific information that seems complementary to general model’s capabilities. See Fig. 6 for the detailed results. As a deeper dive, we examine Mamo-Complex, which is dominated by five problem classes. Fig. 7 (left) shows a per-class breakdown of gpt-oss-20b model without and with hints at inference: accuracy increases are broad and most pronounced on types prone to error related to sign conventions (e.g., within flow-conservation-type constraints) and in complex structural constraints (e.g., subtour elimination constraints in TSP). While IndustryOR and OptMATH contain more heterogeneous types, we observe the same consistent pattern of improvement; see Appendix A.7 for details.
Overall comparison. Our results are summarized in Table 1. As can be seen, the model obtained with the complete OptiMind pipeline, that is, gpt-oss-20b model with SFT, error-analysis hints, multi-turn inference and majority voting, attains accuracy that is close to much larger state-of-the-art models and clearly outperforms open-source models of similar size. For example, when comparing against GPT-5 (single turn, no hints), our model is within $1.9\%$ on OptMATH and is $2.3\%$ and +4.2+4.2 on IndustryOR and Mamo-Complex respectively. In addition, we see sizable improvements over the base gpt-oss-20b model, even when it uses majority voting: $1.3\%$, $4.7\%$, and $5.9\%$ on OptMATH, IndustryOR, and Mamo-complex, respectively. Other OSS models fall significantly behind ours; for example, comparing with SIRL-Gurobi-32B, we obtain improvements of $17\%$$28\%$ across the different benchmarks.
Figure 6: Single-turn accuracy of various models on three cleaned benchmarks w/ and w/o hints. We show consistent improvement in accuracy from adding error-analysis hints.
Figure 7: Left: Accuracy by problem type on Mamo-Complex for gpt-oss-20b (single-turn, no majority voting). Infrequent problem types are omitted. Right: Weighted mean accuracy over five self-debugging turns on the three benchmarks (each weighted by its sample count).
5.3 Additional Ablation Studies
To understand better where the gains come we perform ablation studies on the two main components of our framework – data cleaning and inference strategies.
Ablation on training data. To assess whether our data cleaning improves training, we train three SFT gpt-oss-20b models using progressively higher data quality. (1) Original data: the original OptMATH set plus COPTPY-to-Python/GurobiPy translations from o4-mini, retaining instances whose final solutions match the provided answers; (2) Answer-aligned data: we keep the original problem and final answer, generate 64 solutions with gpt-oss-20b, and keep the sample that matches the original answer even if the answer is imperfect. (3) Cleaned data (ours): we apply our full cleaning pipeline described by §4.3. Table 3 compares single-turn, no-hint accuracy across the three checkpoints. Performance on the original data is poor and, likely due to distribution mismatch in reasoning and lack of reliable thought trajectories. Training on answer-aligned data helps but remains below our cleaned set, suggesting that matching potentially incorrect answers reinforces systematic mistakes and brittle patterns.
Table 3: Single-turn accuracies of SFT variants on different data quality. Numbers are reported by percentage points.
SFT Dataset Hints IndustryOR Mamo-Complex OptMATH
Original Data no $44.48\pm 4.81$ $3.90\pm 2.41 55$$5.31\pm 2.27$
yes $43.36\pm 2.70$ $61.23\pm 2.19$ $80.56\pm 3.99$
Answer-Aligned Data no $66.90\pm 2.62$ $67.75\pm 2.08$ $80.35\pm 1.30$
yes $68.58\pm 2.15$ $84.55\pm 1.90$ $80.24\pm 1.67$
Cleaned Data (ours) no $68.36\pm 2.92$ $74.42\pm 1.15$ $80.35\pm 3.10$
yes $70.26\pm 1.90$ $85.85\pm 0.76$ $81.13\pm 1.59$
Ablation on multi-turn inference. Fig. 7 (right) plots the weighted average accuracy over the number of turns for the base gpt-oss-20b and our SFT model, under the setting of with and without hints. As expected, the accuracy increases monotonically with more turns for both models; however there are diminishing returns that may be taken into account under compute/time vs. accuracy tradeoffs. Overall, the combination SFT + hints is the most effective across benchmarks, especially for small number of turns.
Ablation on Majority Voting (MV). We evaluate the utility of MV when multi-turn correction with tool feedback is available. We perform this ablation on the base gpt-oss-20b model, with and without multi-turn inference. Table 4 shows the effect of MV for the different benchmarks. The main takeaways are: (i) MV and multi-turn are complementary, as both yield sizable gains; (ii) the benefits of MV considerably diminish when multiple turns are applied.
Table 4: Accuracy of gpt-oss-20b across majority voting $K$ and turn counts; hints provided in the prompt. Numbers are reported in percentage points.
IndustryOR Mamo-Complex OptMATH
K 1 turn 5 turns 1 turn 5 turns 1 turn 5 turns
1 72.45{72.45} 76.83{76.83} $83.81$ $89.90$ $79.84$ $86.56$
8 80.10{80.10} $79.69$ $90.04$ $90.72$ 88.36{88.36} 89.14{89.14}
6 Conclusion
We present OptiMind – a framework for formulating mixed-integer linear optimization problems with LLMs, combining structured data cleaning with error-aware prompting, self-consistency, and tool-driven multi-turn correction. Supervised fine-tuning on cleaned data aligns the model with optimization primitives, while targeted hints improve reliability at inference time. The resulting pipeline attains strong performance across multiple benchmarks. While future LLMs may embed more expert knowledge, we believe the principles and techniques of our framework will remain essential. They can be applied to domains such as supply chain management or adapted to enterprise-specific scenarios to drive real-world impact.
Appendix A Appendix
A.1 Sample Problem Description and Output Optimization Model
Below we have a sample problem description from the dataset OptMATH:
Question: Maximize the total profit by determining the optimal production, inventory, and sales plan for three products (Product 0, Product 1, and Product 2) over a five-period planning horizon. The profit per unit sold is \$269 for Product 0, \$282 for Product 1, and \$241 for Product 2. The holding cost for each unit stored is \$15 per period. At the start of the planning horizon (Period 0), the production of each product must equal the sum of its sales and inventory for that period. For each subsequent period, the inventory carried over from the previous period plus the production in the current period must equal the sum of sales and inventory for the current period. The total production time required for all products on each machine type must not exceed the available capacity. The capacities are 480 hours per period for grinders, 320 hours per period for drills, and 160 hours per period for borers. Each unit of Product 0 requires 1 hour on grinders, 1 hour on drills, and 2 hours on borers. Each unit of Product 1 requires 1 hour on grinders, 1 hour on drills, and 2 hours on borers. Each unit of Product 2 requires 1 hour on grinders, 2 hours on drills, and 1 hour on borers. The inventory of each product at the end of any period cannot exceed 80 units. The maximum number of units that can be sold for each product in each period is as follows: Product 0: 48 in Period 0, 43 in Period 1, 58 in Period 2, 58 in Period 3, and 61 in Period 4; Product 1: 54 in Period 0, 56 in Period 1, 45 in Period 2, 46 in Period 3, and 40 in Period 4; Product 2: 57 in Period 0, 52 in Period 1, 68 in Period 2, 40 in Period 3, and 60 in Period 4. At the end of the planning horizon (Period 4), the inventory of each product must be exactly 20 units. Ensure that the final inventory levels meet the target requirements and that production does not exceed machine capacities. # Note: - The Code must include:“‘python import gurobipy as gp from gurobipy import GRB “‘ - Make sure the model variable is named ‘model‘. - Avoid using ”<<>>” in Gurobi constraints; instead, use ”<=<=” or ”>=>=” as appropriate. - Carefully determine whether the variable is an integer or a continuous variable.
Below we have the ground-truth output code formulates this problem as an MILP; the output also includes an initial description of the mathematical formulation in markdown format, which is omited for brevity:
import gurobipy as gp
from gurobipy import GRB
[…]
products = [0, 1, 2]
periods = [0, 1, 2, 3, 4]
machines = [’grinders’, drills’, borers’]
[…]
holding_cost = 15
[…]
grinders’: 480,
drills’: 320,
borers’: 160
[…]
grinders’: {0: 1, 1: 1, 2: 1},
drills’: {0: 1, 1: 1, 2: 2},
borers’: {0: 2, 1: 2, 2: 1}
[…]
0: {0: 48, 1: 43, 2: 58, 3: 58, 4: 61},
1: {0: 54, 1: 56, 2: 45, 3: 46, 4: 40},
2: {0: 57, 1: 52, 2: 68, 3: 40, 4: 60}
[…]
model = gp.Model(”Production_Planning”)
[…]
x = model.addVars(products, periods, name=”Production”)
s = model.addVars(products, periods, name=”Sales”)
I = model.addVars(products, periods, name=”Inventory”)
[…]
model.setObjective(
gp.quicksum(profit[i] * s[i, t] for i in products for t in periods) -
gp.quicksum(holding_cost * I[i, t] for i in products for t in periods),
GRB.MAXIMIZE
[…]
# Initial Inventory Balance (Period 0)
model.addConstrs((x[i, 0] == s[i, 0] + I[i, 0] for i in products), name=”Initial_Balance”)
[…]
model.addConstrs((I[i, t-1] + x[i, t] == s[i, t] + I[i, t] for i in products for t in periods if t >= 1), name=”Inventory_Balance”)
[…]
model.addConstrs((gp.quicksum(machine_time[m][i] * x[i, t] for i in products) <= machine_capacity[m] for m in machines for t in periods), name=”Machine_Capacity”)
[…]
model.addConstrs((s[i, t] <= max_sales[i][t] for i in products for t in periods), name=”Sales_Constraint”)
[…]
model.addConstrs((I[i, t] <= 80 for i in products for t in periods), name=”Inventory_Constraint”)
[…]
model.addConstrs((I[i, 4] == 20 for i in products), name=”Final_Inventory_Target”)
[…]
model.optimize()
[…]
if model.status == GRB.OPTIMAL:
print(”Optimal Solution Found!”)
print(fTotal Profit: \${model.ObjVal:.2f}”)
[…]
for t in periods:
print(fPeriod {t}: Production={x[i,t].X:.2f}, Sales={s[i,t].X:.2f}, Inventory={I[i,t].X:.2f}”)
else:
print(”No optimal solution found.”)
A.2 Problem Classification
For problem classification, we use all 49 seed classes from the OptMATH dataset, finding that these categories are diverse and cover a comprehensive proportion of classes in our training set. Furthermore, we obtain corresponding natural language examples for each problem category from the OptMATH repository. Lu et al. (2025); Table 5 contains the 49 seed class names, and Table 6 shows examples of corresponding examples. When prompting the LLM to classify the problem, we provide all problem categories and natural language examples to assist the language model.
Problem Classification Prompt You are a helpful Assistant with expertise in mathematical modeling and the Gurobi solver. You will be given a natural language question. Please classify the question into one or more of the following categories:
{list of categories}
If the question does not fit any of the above categories, please provide a label for the Category that best describes the problem type.
Category examples:
{natual language example of each category}
Please classify the following question into one or more categories:
{question}
Return only the categories in a python list, without any additional text.
Aircraft Assignment Aircraft Landing Bin Packing
Blending Problem Capacitated Facility Location Problem Capacitated Lot-sizing Problem (CLSP)
Car Selection Assignment Contract Allocation Diet Problem
Factory Planning Problem Flow Shop Scheduling Job Shop
Knapsack Multicommodity Capacitated Network Design MarketShare
Set Multi-Cover PortfolioOptimization Revenue Management Problem
Assignment Problem Set Cover Discrete Lot-sizing and Scheduling Problem
Static Line Planning Structure Based Assignment Problem SupplyChain
TravelingSalesman Facility Dispersion Problem Military Personnel Deployment Problem
Production Planning Problem Facility Dispersion Problem Network Optimization
Lot-Sizing Problem Operations Optimization Capacitated Vehicle Routing Problem with Time Windows (CVRPTW)
Facility Location Problem Cutting Stock Problem Unit Commitment Problem
Farm Planning Transportation, Airline Industry, Resource Allocation Multi-Commodity Transportation Problem
Minimum Cost Flow Problem Assignment Problem Multi-Commodity Network Flow Problem
Transportation Problem Profit Maximization Problem Revenue Maximization Problem
Facility Location Problem Production Planning Problem Team Formulation Problem
Transportation Problem
Table 5: Problem categories obtained from OptMATH
Subclass Natural Language Example
Knapsack A hiker is preparing for a 3-day outdoor hiking trip. They need to select the most valuable combination of equipment and supplies from many available options within the limited backpack capacity. The items include:
* Tent: weight 5kg, value 90 points
* Sleeping bag: weight 3kg, value 80 points* Portable stove: weight 1kg, value 60 points
* Drinking water: weight 2kg, value 70 points
* Dry food: weight 1.5kg, value 65 points
* First aid kit: weight 0.5kg, value 50 points
Assuming the backpack has a maximum weight capacity of 10kg, the hiker’s goal is to select the combination of items with the highest total value while not exceeding the weight limit. Each item must be either taken in its entirety or left behind; partial items cannot be taken.
Team Formulation Problem As a team leader, you have several projects requiring different expertise levels in different areas. You want to assign a set of people to these projects. Each person has a level of expertise in each of the areas. How to assign people to projects to minimize the maximum skill shortage from the project expertise requirement?
Transportation Problem A company has three warehouses (origins) with supplies of 100, 150, and 200 units, respectively. There are four retail stores (destinations) with demands of 80, 120, 90, and 160 units, respectively. The cost of transporting one unit from each warehouse to each retail store is given. The goal is to determine the optimal transportation plan that minimizes the total transportation cost while meeting all supply and demand constraints.
Table 6: Examples of OptMATH classes and their corresponding natural language examples. Full data can be found in https://github.com/optsuite/OptMATH/tree/main/data/generators
A.3 Prompts for Automatically Repairing Missing Data
We design an automatic procedure to fill in missing data in the public available training datasets ORLM Tang et al. (2024) and OptMATH Lu et al. (2025). First, we generate a gurobipy code to each incomplete question by prompting OpenAI’s reasoning models (o1, o3, o4-mini), which often attempts to fill in synthetic data to complete the question. We then extract the synthesized data from the code which corresponds to the missing values. We then use the following prompt to ask o4-mini to produce a modified question with the missing data filled in, where we provide both the original question and the gurobipy code as inputs. Lastly, we manually inspect a subset of 100 repaired questions from OptMATH to make sure that the missing data are successfully filled in.
Prompt for Missing Data Infilling You are given a optimization question and its mathematical solution as well as python code. The question may have missing data or parameters that are filled in by the solution. Your task is to identify if the solution introduces new numbers/values that should be added into the question. Question: {question}
Solution: {completion}
If the question has missing data and the solution provides specific numbers that fills these missing parameters, return the modified question with the missing parameters imputed. Do not modify anything else except for adding in missing data. Otherwise, return NO MISSING DATA. Only return the modified question or NO MISSING DATA, nothing else.
Below is an example question from OptMATH that has missing data infilled by our prompting. The red texts are from the original question with missing data. The green texts are the infilled texts.
Example of Automatic Missing Data Infilling
# Question: Aircraft Assignment for Route Coverage
You are responsible for managing the assignment of aircraft to various
routes for an airline company. The goal is to minimize the total
operational costs while ensuring that all route demands are met
and that the availability of each aircraft type is not exceeded.
#### Aircraft and Routes:
- There are 6 types of aircraft (aircraft_0 to aircraft_5) available
for assignment.
- There are 7 routes (route_0 to route_6) that need to be serviced.
#### Aircraft Availability:
Each aircraft type has a limited number of units available for
assignment:
- Aircraft_0: 4 units available
- Aircraft_1: 5 units available
- Aircraft_2: 4 units available
- Aircraft_3: 4 units available
- Aircraft_4: 5 units available
- Aircraft_5: 5 units available
#### Route Demands:
Each route has a specific demand that must be satisfied by the combined
capabilities of the assigned aircraft. The demands are as follows:
- Route_0: Requires at least 275 units of capacity
- Route_1: Requires at least 213 units of capacity
- Route_2: Requires at least 228 units of capacity
- Route_3: Requires at least 265 units of capacity
- Route_4: Requires at least 226 units of capacity
- Route_5: Requires at least 277 units of capacity
- Route_6: Requires at least 246 units of capacity
#### Aircraft Capabilities: Each aircraft type contributes differently to the capacity of each route. For example: - **Aircraft_0** contributes 100 units to Route_0, 101 units to Route_1, 88 units to Route_2, and so on. - **Aircraft_1** contributes 80 units to Route_0, 88 units to Route_1, 111 units to Route_2, and so on. - Similar contributions are defined for all other aircraft types across all routes. #### Aircraft Capabilities: Each aircraft type contributes differently to the capacity of each route. The capacity contributions $p\_{ij}$ (aircraft $i$ to route $j$) are: - Aircraft_0: [100, 101, 88, 95, 110, 120, 105] - Aircraft_1: [80, 88, 111, 90, 85, 95, 100] - Aircraft_2: [120, 110, 105, 115, 125, 130, 110] - Aircraft_3: [90, 95, 100, 105, 110, 115, 120] - Aircraft_4: [110, 120, 115, 125, 130, 135, 140] - Aircraft_5: [95, 100, 105, 110, 115, 120, 125] #### Operational Costs: Assigning an aircraft to a route incurs a specific cost. For example: - Assigning **Aircraft_0** to **Route_0** costs 3778 units. - Assigning **Aircraft_0** to **Route_1** costs 3344 units. - Assigning **Aircraft_1** to **Route_0** costs 3660 units. - Similar costs are defined for all other aircraft-route combinations. #### Operational Costs: Assigning an aircraft to a route incurs a specific cost $c\_{ij}$ (aircraft $i$ to route $j$): - Aircraft_0: [3778, 3344, 3555, 3666, 3777, 3888, 3999] - Aircraft_1: [3660, 3444, 3555, 3666, 3777, 3888, 3999] - Aircraft_2: [4000, 3888, 3777, 3666, 3555, 3444, $33$] - Aircraft_3: [3555, 3666, 3777, 3888, 3999, 4000, 4111] - Aircraft_4: [3888, 3777, 3666, 3555, 3444, $33$, 3222] - Aircraft_5: [3666, 3555, 3444, $33$, 3222, 3111, 3000]
#### Objective:
Determine the number of each aircraft type to assign to each route
such that:
1. The total operational cost is minimized.
2. The total capacity provided by the assigned aircraft meets
or exceeds the demand for each route.
3. The number of aircraft assigned does not exceed the availability of
any aircraft type.
#### Constraints:
- The total number of aircraft of each type assigned across all routes
must not exceed its availability.
- The combined capacity of all aircraft assigned to a route must meet
or exceed the route’s demand.
- The number of aircraft assigned to any route must be a non-negative
integer.
A.4 Multi-Turn Inference Prompts
Algorithm 1 MultiTurnInferencewithMajorityVoting
problem instance $Q$, error-analysis library $H$ (maps problem type $\to$ list of (error, hint) pairs), LLM Generator $G$, majority voting number $K$, number of correction rounds $M$.
$t\leftarrow\textsc{ClassifyType}_{G}(Q)$ $\triangleright$ predict problem type
$\texttt{hints}\leftarrow H[t]$ $\triangleright$ retrieve hints of corresponding type
$\big(s_{1}^{(0)},\cdots s_{K}^{(0)}\big)\leftarrow\textsc{FirstTurnGenerate}_{G}(Q,\texttt{hints},K)$ $\triangleright$ generate $K$ solutions
$a^{(0)},\texttt{stdout}^{(0)},\texttt{stderr}^{(0)}\leftarrow\textsc{GetMajorityResults}\big(s_{1}^{(0)},\cdots s_{K}^{(0)}\big)$
for $m=1$ to M1M-1 do
  $\big(s_{1}^{(m)},\cdots s_{K}^{(m)}\big)\leftarrow\textsc{SelfCorrectionGenerate}_{G}(Q,\texttt{stdout}^{(m-1)},\texttt{stderr}^{(m-1)},K)$ $\triangleright$ get $K$ self-correction responses
  $a^{(m)},\texttt{stdout}^{(m)},\texttt{stderr}^{(m)}\leftarrow\textsc{GetMajorityResults}\big(s_{1}^{(m)},\cdots s_{K}^{(m)}\big)$
end for
return $\hat{a}^{(M-1)}$ $\triangleright$ Return the final result
Algorithm 2 GetMajorityResults($s_{1},\cdots s_{K}$)
Solution trajectories $s_{1},\cdots s_{K}$
for $k=1$ to $K$ do
  $\big(a_{k},\texttt{stdout}_{k},\texttt{stderr}_{k}\big)\leftarrow\textsc{ExtractAnswer}(s_{k})$ $\triangleright$ Extract code and get system output for each solution
end for
$\hat{k}\leftarrow\textsc{GetMajorityVoteIndex}\big(a_{1},\cdots,a_{K}\big)$
[…]
We use the following prompts during our multi-turn inference. At the first turn, we prompt the LLM with the question and asks for the corresponding gurobipy code. Optionally, we may include an error-analysis hint for the problem’s category, which we find generally improve the performance. Algorithm 1 shows a pseudocode of our inference strategy.
First Turn User Prompt (Not Using Error Analysis Hint) You are an expert in optimization and mixed integer programming. You are given an optimization problem and you need to solve it using gurobipy.
{question}
Reason step by step before generating the gurobipy code. When you respond, first think carefully. After thinking, output the math modeling of the problem. Finally output a ‘‘‘python ...‘‘‘ code block that solves the problem.
The code must include: import gurobipy as gp
from gurobipy import GRB
First Turn User Prompt (Using Error Analysis Hint) You are an expert in optimization and mixed integer programming. You are given an optimization problem and you need to solve it using gurobipy.
{question}
Below are hints for avoiding common mistakes often seen for this problem type. Avoid them if applicable.
Error analysis #1: {Error Summary 1}, {Hint 1} ... Error analysis #k: {Error Summary k}, {Hint k}
Instructions for applying error-analysis hints:
- Review the provided hints and identify which ones are applicable to this problem.
- Please apply ALL applicable hints.
- Before applying any hint or writing constraints, check the sign and direction of every variable and coefficient for consistency (e.g., profit = revenue - cost; capacities as <= flow conservation as =).
General modeling checklist (follow rigorously):
- Units: Use correct units everywhere and ensure the objective’s units match the goal (e.g., dollars for cost/profit, distance for TSP, time for scheduling). Do not mix units (e.g., minutes with hours, dollars with 1000 dollars) without converting.
Reason step by step before generating the gurobipy code. When you respond, first think carefully. After thinking, output the math modeling of the problem. Finally output a ‘‘‘python ...‘‘‘ code block that solves the problem.
The code must include: import gurobipy as gp
from gurobipy import GRB
From the LLM’s response, we extract and execute the gurobipy code to obtain standard output (STDOUT) and standard error (STDERR). Our multi-turn pipeline then feeds this execution feedback back to the LLM, prompting it to self-correct its previous solution, as shown below.
Self-Correction Feedback Prompt Running the given gurobipy code block result in the following standard output and error.
Please analyze the standard output and error and provide your thoughts on the correctness of the code and the output, and if it is not correct, provide a corrected version of the code.
The standard output and error message (if any) for your generated code are as follows:
[STDOUT] {Standard Output of the gurobipy program (gurobipy log)}
[STDERR] {Standard Error of the gurobipy program (if exists)}
Based on the output and error message (if any), please think step by step in the <think> ... </think> block to determine if the code is correct.
[…]
# Note: - The Code must include:‘‘‘python import gurobipy as gp from gurobipy import GRB ‘‘‘
- Make sure the model variable is named ‘model‘.
- Avoid using "<" and ">" in Gurobi constraints; instead, use "<=" or ">=".
- Carefully determine whether the variable is an integer or continuous.
A.5 Example of the Effect of Multi-Turn Inference
We describe an example where multi-turn inference corrected the formulation.
This instance is production optimization problem from the IndustryOR dataset. In this instance, a factory has a warehouse for holding inventory with capacity of 15,000 cubic meters. In addition, it can use an external warehouse (at an additional cost).
The first turn of the inference produced a model with the incorrect constraint e[t] <= vol - wh_vol_cap, where e[t] is a decision variable choosing how many cubic meters of external warehouse to use in period $t$, vol denotes the total volume of (internal + external) warehouse used, and wh_vol_cap =15,000=15,000 is the internal warehouse capacity. Notice that this constraint is incorrect because it allows the volume used vol to be bigger than the capacity wh_vol_cap while having external warehouse usage e[t] =0=0. The correct inequality is in the opposite direction, namely “ e[t] >= vol - wh_vol_cap”, which forces that the external warehouse usage to cover for any excess over the internal warehouse capacity.
The second turn of the inference used the output from the first turn (in particular, the first-turn formulation and the optimal decisions of this first model) to detect this mistake and come up with the correct constraint. Here is the relevant snipped from the second-turn reasoning:
Thought of the second turn of inference (snippet) Volume per month: compute volume each month. July: 0.2*45k + 0.4*15k = 9000 + 6000 = 15000 cubic meters. That equals capacity. So no external. August: 0.2*75k + 0.4*0 = 15000. OK. September: 0.2*127.5k + 0.4*0 = 25500. Exceeds 15000 by 10500. But model forced e_t <= vol - wh_vol_cap. So e_t would be 10500. But output shows e_t 0. So maybe e_t not used? Let’s check: The constraint e[t] <= vol - wh_vol_cap. For September vol=25500, wh_vol_cap=15000 => RHS=10500. So e_t <= 10500. But objective includes 0.5*e_t. To minimize cost, e_t should be 0. But we didn’t enforce e_t >= vol - wh_vol_cap? No, we only enforced e_t <= vol - wh_vol_cap. That allows e_t to be 0 even if vol > wh_vol_cap. But we need to enforce that external volume equals excess volume. We used inequality incorrectly. We should enforce e_t >= vol - wh_vol_cap, not <=.
A.6 Experiment Details
A.6.1 Training and Evaluation Setups
We performed Supervised fine-tuning (SFT) on a single node with eight NVIDIA HGX B200 GPUs. All solution generation in the data-cleaning pipeline and evaluation tasks were run on 4 compute nodes, each with eight 80GB NVIDIA H100 GPUs. We use gpt-oss-20b as the base model and adapt the Verl framework for SFT. Concretely, we train from the unsloth/openai-gpt-oss-20b-BFloat16 variant to avoid MXFP4 precision pitfalls and to leverage Unsloth’s compatibility fixes for gpt-oss, and uses FSDP2 strategy to distribute training across the eight GPUs. Key SFT hyperparameters and optimizer settings are listed in Table 7.
Table 7: Generation / evaluation sampling parameters.
Parameter Value
Batch size 16
Learning rate 1e-6
Max prompt length 16384
Max response length 16384
Prompt length truncation left
Max epochs 3
For solution generation in data cleaning, we host openai/gpt-oss-20b with vLLM. For evaluation, we serve all SFT checkpoints and the BF16 base (unsloth/openai-gpt-oss-20b-BFloat16) via SGLang (vLLM did not support our Unsloth-adapted variant at the time when we conduct the evaluation). We use top-p decoding for all generation and inference; the common sampling settings are summarized in Table 8.
Table 8: Generation / evaluation sampling parameters.
Parameter Value
Temperature 0.6
Top p 0.95
Max tokens 60000
A.7 Details of Error-Analysis Hints at Inference
Figure 9 presents a per-problem-class breakdown of the impact of using error-analysis and associated hints at inference on the gpt-oss-20b model.
Figure 8: Accuracy by problem type on IndustryOR (top) and OPTMath (bottom), measured via single-turn evaluation with gpt-oss-20b (no majority voting). Problem types occurring in less than 2.5% of instances are omitted in the figure.
Figure 9: Number of error-hint pairs that experts collected per class.
A.8 Details of Training Data Cleaning
Below we provide an example of cleaning training instance of the Job Shop problem with missing data. We describe the semi-automated procedure for infilling missing data in Appendix A.3.
Job Shop Problem (Before) # Question:In a manufacturing facility, six distinct jobs must be processed on a set of machines. Each job consists of two operations, and each operation must be performed on a specific machine. The goal is to schedule these operations to minimize the total completion time, known as the makespan. Each operation has a processing time of 1 unit, and operations within the same job must follow a specific sequence, with at least 1 unit of time between consecutive operations. Each machine can process only one operation at a time, and the order of operations on shared machines is determined by binary decision variables. These binary variables ensure that no two operations overlap on the same machine. The makespan must be at least as large as the completion time of every operation, and all start times must be non-negative. The binary variables take values of 0 or 1, and a large constant (100,000) is used in the machine capacity constraints to enforce the order of operations. The objective is to determine the start times of all operations and the order in which they are processed on each machine, ensuring that all constraints are satisfied and the makespan is minimized.
# Note:
- The Code must include:‘‘‘python
import gurobipy as gp
from gurobipy import GRB
- Make sure the model variable is named ‘model‘.
- Avoid using "<<" and ">>" in Gurobi constraints; instead, use "<=<=" or ">=>=" as appropriate.
- Carefully determine whether the variable is an integer or a continuous variable.
Job Shop Problem (After) # Question:In a manufacturing facility, six distinct jobs must be processed on a set of machines. Each job consists of two operations, and each operation must be performed on a specific machine. In particular, for Job 0: Operation 0 is assigned to Machine 2, and Operation 1 to Machine 1, for Job 1: Operation 0 is assigned to Machine 0, and Operation 1 to Machine 2, for Job 2: Operation 0 is assigned to Machine 1, and Operation 1 to Machine 0, for Job 3: Operation 0 is assigned to Machine 2, and Operation 1 to Machine 2, for Job 4: Operation 0 is assigned to Machine 1, and Operation 1 to Machine 0, and for Job 5: Operation 0 is assigned to Machine 0, and Operation 1 to Machine 1. The goal is to schedule these operations to minimize the total completion time, known as the makespan. Each operation has a processing time of 1 unit, and operations within the same job must follow a specific sequence, with at least 1 unit of time between consecutive operations. Each machine can process only one operation at a time, and the order of operations on shared machines is determined by binary decision variables. These binary variables ensure that no two operations overlap on the same machine. The makespan must be at least as large as the completion time of every operation, and all start times must be non-negative. The binary variables take values of 0 or 1, and a large constant (1000) is used in the machine capacity constraints to enforce the order of operations. The objective is to determine the start times of all operations and the order in which they are processed on each machine, ensuring that all constraints are satisfied and the makespan is minimized.
# Note:
- The Code must include:‘‘‘python
import gurobipy as gp
from gurobipy import GRB
- Make sure the model variable is named ‘model‘.
- Avoid using "<<" and ">>" in Gurobi constraints; instead, use "<=<=" or ">=>=" as appropriate.
- Carefully determine whether the variable is an integer or a continuous variable.
A.9 Details of Benchmarks Cleaning
This section dives into some of the issues we observed in the benchmarks data. These fall under six categories: missing data, ambiguous problem descriptions, integral vs fractional variables, wrong solutions, infeasible problems, and out-of-scope problems. We next provide additional details specific examples that highlight these challenges.
A.9.1 Missing Data
We observe that many problems contain missing values for key parameters. For example, an assignment problem in OptMATH reads “For example, assigning aircraft_0 to route_0 costs 2552 units, to route_1 costs 4340 units, and so on” without providing the additional costs. Interestingly, when data is missing, the OpenAI o-series models synthesize reasonable values for the missing parameters during the formulation generation process. We therefore manually identify problems with missing data, and leverage the fabricated values in their solutions to fill them back into the question description, followed by manual inspection to validate correctness.
A.9.2 Ambiguous Problem Descriptions
We also observe that many problems contain ambiguities or inconsistencies. For example, a facility location problem in IndustryOR had the following objective function “to achieve the goal of minimizing costs and maximizing coverage area”, thus containing two conflicting objectives. These issues are corrected via manual inspection.
A.9.3 Integral vs Fractional Variables
Many problems exhibit ambiguity regarding whether decision variables are integer or continuous. For example, a production optimization problem in IndustryOR refers to “units of products”, while the reported ground truth corresponds to the production of fractional products. We resolve this by either updating the ground truth value, adding explicit clarification sentences that enforce integrality, or when both interpretations are reasonable, by providing two ground-truth answers corresponding to the integer and fractional formulations.
Example of Integer vs. Fractional A company produces product A and product B. Each unit of product A sold generates a profit of £30, while each unit of product B sold generates a profit of £10. The company can allocate a maximum of 40 hours per week for production. Producing one unit of product A requires 6 hours, while producing one unit of product B requires 3 hours. Market demand requires that the quantity of product B produced must be at least three times the quantity of product A. The storage space occupied by product A is four times that of product B, and a maximum of four units of product A can be stored per week.
Formulate a model for this problem.
Reported ground truth: 146.667
Corrected ground truth: 160
A.9.4 Wrong Solutions
The benchmarks also contained problems where the provided ground truth value was incorrect. For instance, some minimum-cost flow problems in the OptMATH dataset reported wrong optimal costs (e.g., one problem reported optimal cost of 6 as the ground truth, while the correct optimal value was 4). These errors were all corrected with manual inspection by optimization experts.
Example of Wrong Solution You are responsible for managing the distribution of emergency medical supplies across eight cities in a region. Each city has a specific supply of medical resources and a demand that must be met to ensure adequate healthcare coverage. The goal is to minimize the total transportation cost while ensuring that all cities receive the necessary supplies and that no distribution routes exceed their capacity.
City Supply and Demand:
- **City 0** has a net demand of 1 unit of medical supplies.
- **City 1** has a balanced supply and demand (net demand of 0 units).
- **City 2** has a net supply of 1 unit of medical supplies.
- **City 3** has a net demand of 2 units of medical supplies.
- **City 4** has a balanced supply and demand (net demand of 0 units).
- **City 5** has a balanced supply and demand (net demand of 0 units).
- **City 6** has a net supply of 2 units of medical supplies.
- **City 7** has a balanced supply and demand (net demand of 0 units).
Transportation Costs: The cost of transporting medical supplies between cities varies depending on the route. Below are the transportation costs per unit of supplies:
- **From City 0**: To City 1 costs 3, to City 2 costs 2, to City 3 costs 2, to City 4 costs 2, to City 5 costs 3, to City 6 costs 3, and to City 7 costs 1. - **From City 1**: To City 0 costs 1, to City 2 costs 2, to City 3 costs 3, to City 4 costs 1, to City 5 costs 2, to City 6 costs 1, and to City 7 costs 2. - **From City 2**: To City 0 costs 2, to City 1 costs 2, to City 3 costs 3, to City 4 costs 3, to City 5 costs 2, to City 6 costs 1, and to City 7 costs 2. - **From City 3**: To City 0 costs 1, to City 1 costs 2, to City 2 costs 1, to City 4 costs 3, to City 5 costs 3, to City 6 costs 2, and to City 7 costs 3. - **From City 4**: To City 0 costs 3, to City 1 costs 2, to City 2 costs 1, to City 3 costs 1, to City 5 costs 3, to City 6 costs 2, and to City 7 costs 2. - **From City 5**: To City 0 costs 1, to City 1 costs 2, to City 2 costs 1, to City 3 costs 2, to City 4 costs 1, to City 6 costs 2, and to City 7 costs 1. - **From City 6**: To City 0 costs 2, to City 1 costs 3, to City 2 costs 1, to City 3 costs 1, to City 4 costs 1, to City 5 costs 1, and to City 7 costs 1. - **From City 7**: To City 0 costs 1, to City 1 costs 1, to City 2 costs 3, to City 3 costs 1, to City 4 costs 2, to City 5 costs 3, and to City 6 costs 2.
Route Capacity Constraints: Each route between cities has a maximum capacity for transporting medical supplies:
- **From City 0**: To City 1 (7 units), to City 2 (7 units), to City 3 (7 units), to City 4 (7 units), to City 5 (8 units), to City 6 (8 units), and to City 7 (8 units). - **From City 1**: To City 0 (8 units), to City 2 (7 units), to City 3 (8 units), to City 4 (8 units), to City 5 (7 units), to City 6 (7 units), and to City 7 (9 units). - **From City 2**: To City 0 (8 units), to City 1 (7 units), to City 3 (7 units), to City 4 (7 units), to City 5 (7 units), to City 6 (9 units), and to City 7 (7 units). - **From City 3**: To City 0 (7 units), to City 1 (7 units), to City 2 (9 units), to City 4 (8 units), to City 5 (7 units), to City 6 (7 units), and to City 7 (9 units). - **From City 4**: To City 0 (9 units), to City 1 (7 units), to City 2 (8 units), to City 3 (9 units), to City 5 (7 units), to City 6 (7 units), and to City 7 (7 units). - **From City 5**: To City 0 (7 units), to City 1 (8 units), to City 2 (9 units), to City 3 (9 units), to City 4 (8 units), to City 6 (9 units), and to City 7 (8 units). - **From City 6**: To City 0 (9 units), to City 1 (8 units), to City 2 (7 units), to City 3 (8 units), to City 4 (8 units), to City 5 (7 units), and to City 7 (8 units). - **From City 7**: To City 0 (9 units), to City 1 (8 units), to City 2 (7 units), to City 3 (9 units), to City 4 (9 units), to City 5 (8 units), and to City 6 (8 units).
City Capacity Constraints: Each city has a maximum capacity for receiving medical supplies:
- **City 0**: Can receive up to 19 units.
- **City 1**: Can receive up to 15 units.
- **City 2**: Can receive up to 15 units.
- **City 3**: Can receive up to 14 units.
- **City 4**: Can receive up to 15 units.
- **City 5**: Can receive up to 15 units.
- **City 6**: Can receive up to 14 units.
- **City 7**: Can receive up to 16 units.
Objective: Your task is to determine the optimal distribution of medical supplies between the cities to minimize the total transportation cost while ensuring that all cities meet their supply and demand requirements, no route exceeds its capacity, and no city exceeds its receiving capacity.
Reported ground truth: 6.
Correct ground truth: 4.
A.9.5 Infeasible Problems
Beyond problems with wrong solutions, we also observe problems that are infeasible. An example is provided below. We fix these problems by appropriately updating the data so that the problem admits a feasible solution.
Example of Infeasible Problem "A university computer lab hires 4 undergraduates (designated 1, 2, 3, and 4) and 2 graduate students (designated 5 and 6) for duty answering questions. The maximum duty hours from Monday to Friday and the hourly wage for each person are shown in Table 5-9.
Table 5-9
Student ID,Wage (CNY/h),Monday,Tuesday,Wednesday,Thursday,Friday
1,10.0,6,0,6,0,7
2,10.0,0,6,0,6,0
3,9.9,4,8,3,0,5
4,9.8,5,5,6,0,4
5,10.8,3,0,4,8,0
6,11.3,0,6,0,6,3
The lab operates from 8:$0$ AM to 10:$0$ PM, and there must be one and only one student on duty during open hours. It is also required that each undergraduate must work at least 8 hours per week, and each graduate student must work at least 7 hours per week. Additionally, supplement the following requirements: each student can work no more than 2 shifts per week, and no more than 3 students can be scheduled for duty each day. Based on these conditions, establish a new mathematical model."
Example of Fixed Infeasible Problem "A university computer lab hires 4 undergraduates (designated 1, 2, 3, and 4) and 2 graduate students (designated 5 and 6) for duty answering questions. The maximum duty hours from Monday to Friday and the hourly wage for each person are shown in Table 5-9.
Table 5-9
Student ID | Wage (CNY/h) | Monday | Tuesday | Wednesday | Thursday | Friday
1 | 10.0 | 6 | 0 | 6 | 0 | 7
2 | 10.0 | 0 | 6 | 0 | 6 | 7
3 | 9.9 | 4 | 8 | 4 | 0 | 5
4 | 9.8 | 5 | 5 | 6 | 0 | 4
5 | 10.8 | 4 | 0 | 4 | 8 | 0
6 | 11.3 | 5 | 6 | 0 | 6 | 3
The lab operates from 8:$0$ AM to 10:$0$ PM, and there must be one and only one student on duty during open hours. It is also required that each undergraduate must work at least 8 hours per week, and each graduate student must work at least 7 hours per week. Additionally, each student can work no more than 2 shifts per week, and no more than 3 students can be scheduled for duty each day.
Based on these conditions, establish a mathematical model to determine the work schedule that satisfies all requirements."
A.9.6 Out-of-Scope Problems
Finally, we also observe a small number of non-linear problems. For instance, OptMATH contains certain second-order cone programming and quadratically constrained programming problems. Since our focus in this work is on (mixed integer) linear optimization, we manually omit these non-linear instances from the benchmarks. We also find that the available training datasets contain no comparable nonlinear training data, so including such instances would create a train–test mismatch.
Example of Non-Linear Problem Minimize the objective function with cost coefficients -0.1919592146476727 for x[0], -1.473647303839492 for x[1], and 2.304735407761341 for x[2]. The problem is subject to three strict equality constraints: (1) -1.635895473616174 x[0] - 0.3973211001807447 x[1] + 0.9471007364101932 x[2] must equal -1.151224267166901, labeled as linear_eq[0]; (2) -0.5249535603075616 x[0] + 0.3668073807989349 x[1] - 0.7858336216136411 x[2] must equal 0.2111440590441619, labeled as linear_eq[1]; and (3) -1.91474276776438 x[0] - 0.0172618223950067 x[1] - 0.4534063203075578 x[2] must equal -0.0637133383926026, labeled as linear_eq[2].
Additionally, four auxiliary variables y_0[0], y_0[1], y_0[2], and y_0[3] are introduced, each subject to linear constraints: (1) -0.1884394252535835 x[0] + 0.0177104249784969 x[1] + 0.0986130842848287 x[2] + y_0[0] must equal 0.0043955531724021, labeled as R3; (2) -0.0708818331240574 x[0] - 0.0215819660876361 x[1] - 0.0616654799104094 x[2] + y_0[1] must equal 0.0879019749446116, labeled as R4; (3) -0.0907143884234421 x[0] + 0.0750697315409076 x[1] - 0.1169779589661307 x[2] + y_0[2] must equal 0.0357726532121181, labeled as R5; and (4) -0.1820309484664378 x[0] + 0.0261698752429371 x[1] + 0.0667294286775777 x[2] + y_0[3] must equal 0.0118088378651355, labeled as R6.
A decision variable t_0[0] is introduced, subject to the constraint -0.2396501315769039 x[0] + 0.3255209569765619 x[1] + 0.2624492403208943 x[2] + t_0[0] must equal 1.588464205354944, labeled as R7. A second-order cone constraint labeled as qc0 is imposed, ensuring that the sum of the squares of y_0[0], y_0[1], y_0[2], and y_0[3] does not exceed the square of t_0[0], expressed as - t_0[0]ˆ2 + y_0[0]ˆ2 + y_0[1]ˆ2 + y_0[2]ˆ2 + y_0[3]ˆ2 <= 0. All decision variables, including x[0], x[1], x[2], y_0[0], y_0[1], y_0[2], y_0[3], and t_0[0], are free to take any real value. The goal is to determine the optimal values for these variables to minimize the objective function while satisfying all constraints. This is a Second-Order Cone Programming (SOCP) problem."